OpenStudy (anonymous):

Find Three Ordered Pair Solutions to the Equation: 2y=3x+11

OpenStudy (anonymous):

@ajprincess

OpenStudy (anonymous):

A math question in history:) Was this homework overdue;)

OpenStudy (anonymous):

OOPS Thought It Was Math

OpenStudy (ajprincess):

put 1 in place of x and solve for y. Can u tell me vat u get @DamnImBad?

OpenStudy (anonymous):

Close enough you got the right website;)

OpenStudy (anonymous):

(0,7) (1,7) (2,7)

OpenStudy (ajprincess):

(1,7) is right:) Bt other two aren't right.

OpenStudy (ajprincess):

2y=3x+11 Let us put x=0 2y=3*0+11 2y=0+11 2y=11 y=11/2 (0,11/2) Let us put x=2 2y=3*2+11 2y=6+11 2y=17 y=17/2 (2,17/2) Getting it @DamnImBad?

OpenStudy (anonymous):

Yes

OpenStudy (ajprincess):

ok 5n:)

OpenStudy (anonymous):

Huh ?